OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2?

hartnn (hartnn):

the slope point form of line is y-y1 = m(x-x1) where (x1,y1) is the point(here (2,5)) and m is the slope (here m=2) now can u find the equation of line ?

OpenStudy (anonymous):

y-5=2(x-2) y-5=2x-4 y=2x+1

hartnn (hartnn):

sorry for late reply. thats is correct! y=2x+1 is the answer :)

OpenStudy (anonymous):

:) whoop! i like feeling smart.